OpenStudy (cheska_p):

Find the vertex, focus, directrix, and axis of the parabola: x^2 -4x+3y-5=0 I understand how to find everything the question is asking, but not when the parabola is in that form. How do I get that particular parabola to be in standard form?

OpenStudy (anonymous):

complete the square (hope i don't steer you wrong this time) btw i saw my mistake on the last one, it was \(20+40=60\) i believe do you know how to complete the square?

OpenStudy (anonymous):

\[x^2 -4x+3y-5=0 \] complete the square on the \(x\) term to get \[(x-2)^2+3y-5-4=0\] or \[(x-2)^2+3y-9=0\]

OpenStudy (cheska_p):

so how do we complete the square?

OpenStudy (anonymous):

you have \[x^2-4x\] lets ignore everything else

OpenStudy (cheska_p):

okay so far I'm following

OpenStudy (anonymous):

half of \(-4\) is \(-2\) so you write \[x^2-4x=(x-2)^2-4\]

OpenStudy (anonymous):

the reason being if you square \(x-2\) you get \[x^2-4x+4\]so you have added 4, now you need to subtract it off to make it equal to what you started with

OpenStudy (anonymous):

that is how we ended up with \[(x-2)^2+3y-9=0\]

OpenStudy (anonymous):

now make it \[-3y+9=(x-2)^2\]

OpenStudy (anonymous):

factor out the \(-3\) (carefully) and get \[-3(y-3)=(x-2)^2\] and you can read off the vertex from that one

OpenStudy (anonymous):

you good from there? it should be in the form you want it now

OpenStudy (anonymous):

how did you go from summing a geometric series to this? quite random

OpenStudy (cheska_p):

Yes I'm just writing everything out. Hahaha, I doing a practice test for my test tomorrow so it covers a couple sections

OpenStudy (cheska_p):

I'm*

OpenStudy (anonymous):

i see i hope you don't get a bouncing ball problem if you do, remember to take the distance that it falls outside the series also remember that \[a+ar+ar^2+ar^3+...=\frac{a}{1-r}\]

OpenStudy (anonymous):

in any case you should be done here \[-3(y-3)=(x-2)^2\]gives you the vertex of \((2,3)\)

OpenStudy (cheska_p):

yes I got that :)

OpenStudy (anonymous):

then since \(4p=3\) you get \(p=\frac{3}{4}\) and the directix is \(\frac{3}{4}\) units above the vertex, focus is \(\frac{3}{4}\) units below

OpenStudy (anonymous):

let me make sure that is right so i don't steer you wrong again

OpenStudy (anonymous):

no, it is right directrix \(y=3+\frac{3}{4}=\frac{15}{4}\)

OpenStudy (anonymous):

and since \(3-\frac{3}{4}=\frac{9}{4}\) the focus is \((2,\frac{9}{4})\)

OpenStudy (cheska_p):

Hahaha I was just about to check that with you, thank you. I got them right

OpenStudy (anonymous):

great!